Shift summary for eve., 10/21/99

Jianbo Wang (jwang@physics.arizona.edu)
Thu, 21 Oct 1999 22:03:34 -0700 (MST)


Shift Summary for Eve Shift, 10/21/99

On Shift: RickK, Jianbo
Sydney and Julie have been here for a few hours

Description of shift: Finish Y2K test successfully, low ET threshold test

18:20 Restore of ktev1-4, atev1-3 orginal. Sydney started slow DAQ.

18:35 Modified L3PROC cosmics +BEAM_MUON (used by laser_scan)
Files to use 1 less L3 process on filter planes- this should fix
SHM_SUM error seen at init of LASER_SCAN just before the 3-week
shutdown

18:50 Successful cosmic run!!! (calib trigs only)
End of Y2K test

19:00 Etot+colling turned back on ( need to turn off ETOT & CsI when
finished). Beam triggers didn't work earlier..
B16 (random pulse) started working again while degugging
Beam trigger with L2 still hang, leave it for now.

Creat new config: COSMIC_ETOT

21:17 start COSMIC_ETOT run 14876, cannot see anything in ETOT
Restart run with dusk output to debug monitor code.

21:26 Start run 14877, /ktev3e/SPLIT/cos has the data. While debugging
ET monitor software, found a bug that ginore calib triggers... fix
it, remake ktevmon, and remade L3_driver_mon_cosm

RickK ran cosmic run to test ET low threshold. Everything is fine.

22:01 RK start run 14878 with debugged monitor code for ETOT, unplug
on|-spill cosmic veto at bottom of RR 164 to increase rate.

23:25 Finish gas check, ArEth rack is out of range, it read 13